# Functors on the category of finite sets revisited

### v2

@@ 10.1. Calculating $\mathrm{Hom}(\Lambda^{t}(P^{\mathbf{FA}}),P^{\mathbf{FA}}_{\mathbf{n}})$
Clearly $\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{t})=0$ if $s>t$, so we may assume that $s\leq t$. Now, there is an isomorphism of $\mathbbm{k}(\mathfrak{S}_{t}\times\mathfrak{S}_{s}^{\mathrm{op}})$-modules $\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{t})\cong\mathbbm{k}\mathfrak{S}_{t}/\mathfrak{S}_{t-s}$, where the actions are inherited from the left and right regular actions on $\mathbbm{k}\mathfrak{S}_{t}$. It follows that one has the isomorphism $\mathsf{sgn}_{t}\otimes_{\mathfrak{S}_{t}}\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{t})\cong\mathsf{sgn}_{t}|_{\mathfrak{S}_{t-s}}\otimes_{\mathfrak{S}_{t-s}}\mathsf{triv}_{t-s}$ with respect to the residual $\mathbbm{k}\mathfrak{S}_{s}$-action arising from $\mathsf{sgn}_{t}|_{\mathfrak{S}_{s}}$ on the right hand side. This is clearly zero unless $t-s\in\{0,1\}$ in which case the representation identifies as stated. ∎
@@
This result is related to the short exact sequence (4.7) in Remark 4.10.
@@
$0\rightarrow\mathbbm{k}_{\mathbf{0}}^{\mathbf{FA}^{\mathrm{op}}}\rightarrow\mathbbm{k}^{\mathbf{FA}^{\mathrm{op}}}\rightarrow\overline{\mathbbm{k}}^{\mathbf{FA}^{\mathrm{op}}}\rightarrow 0$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},P^{\mathbf{FA}}_{\mathbf{n}})]=[\mathbbm{k}{\mathbf{FS}}(\mathbf{n},-)]\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(\mathbf{n},\mathbf{k}))\otimes[\mathsf{sgn}_{k-1}].$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},P^{\mathbf{FA}}_{\mathbf{n}})]=[\overline{P^{\mathbf{FA}}_{\mathbf{n}}}]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}P^{\mathbf{FA}}_{n}(\mathbf{k}))\otimes\mathcal{S}(k-1),$
The hypothesis $n>0$ ensures that $\overline{P^{\mathbf{FA}}_{\mathbf{n}}}=P^{\mathbf{FA}}_{\mathbf{n}}$. Thus, using the identity as $\mathbbm{k}\mathbf{FB}$-bimodules $\mathbbm{k}\mathbf{FA}\cong\mathsf{triv}\odot_{\mathbf{FB}}\mathbbm{k}{\mathbf{FS}}\cong\mathbbm{k}{\mathbf{FS}}\odot_{\mathbf{FB}}\mathsf{triv}$ given in Example 2.6, together with Proposition 8.3, one has $[\overline{P^{\mathbf{FA}}_{\mathbf{n}}}]\odot_{\mathbf{FB}}\mathcal{S}(0)=[\mathbbm{k}{\mathbf{FS}}(\mathbf{n},-)]$, which accounts for the first term in the statement.
@@
$\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}P^{\mathbf{FA}}_{n}(\mathbf{k}))\otimes\mathcal{S}(k-1)$ $=$ $\sum_{k\geq 1}\Big{(}M(\mathbf{n},\mathbf{k-1})\otimes\mathcal{S}(k-1)\ \oplus\ M(\mathbf{n},\mathbf{k})\otimes\mathcal{S}(k-1)\Big{)}$
$=$ $\sum_{k\geq 1}M(\mathbf{n},\mathbf{k})\otimes(\mathcal{S}(k-1)\oplus\mathcal{S}(k))$
@@
where the second equality reindexes the sum. The result follows. ∎
@@
Hence, using Proposition 8.3, one can calculate $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]$:
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]=[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}\Big{(}[\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(-,\mathbf{k})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\Big{)}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}},$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]=[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P^{\mathbf{FA}}_{*}})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0).$
The result follows by using the identification of $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P^{\mathbf{FA}}_{*}})]$ given by Theorem 10.6. ∎
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\mathbb{P}_{*})]$ $=$ $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]\quad+\quad\sum_{\ell\geq 0}[\mathsf{sgn}_{\ell}]_{\mathbf{FB}^{\mathrm{op}}}\boxtimes[\mathsf{sgn}_{\ell+1}]_{\mathbf{FB}}$
$=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}\Big{(}[\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(-,\mathbf{k})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\Big{)}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$
@@ 11. Calculating $\mathrm{Hom}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})$
The purpose of this section is to calculate the underlying $\mathbbm{k}\mathbf{FB}$-bimodule of $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})$. Theorem 10.8 described $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]$ in the Grothendieck group of $\mathbbm{k}\mathbf{FB}$-bimodules. We exploit the short exact sequence
$0\rightarrow\Lambda^{\bullet+1}(\overline{P})\rightarrow\mathbb{P}_{\bullet}\rightarrow\overline{P}^{\otimes\bullet}\rightarrow 0$
@@ 11.1. Morphisms out of $\Lambda^{s}(\overline{P})$
using the $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-module structure of Proposition 2.11 on $\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{s})$. Moreover, with respect to this isomorphism, for $s,t\in\mathbb{N}$, the short exact sequence $0\rightarrow\Lambda^{s+1}(\overline{P})\rightarrow\Lambda^{s+1}(P^{\mathbf{FA}})\rightarrow\Lambda^{s}(\overline{P})\rightarrow 0$ induces a short exact sequence
$0\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s+1}(P^{\mathbf{FA}}),P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s+1}(\overline{P}),P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow 0$
@@
The first statement is an immediate consequence of the natural isomorphism $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes s},P^{\mathbf{FA}}_{\bullet})\cong\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{s})$ provided by Theorem 4.8. That result also implies that the inclusion $\Lambda^{s+1}(\overline{P})\hookrightarrow\Lambda^{s+1}(P^{\mathbf{FA}}_{\mathbf{1}})$ induces a surjection
@@ 11.2. Calculating $\mathrm{Hom}(\Lambda^{s}(\overline{P}),\overline{P}^{\otimes*})$
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),\overline{P}^{\otimes\bullet}/\Lambda)\odot_{\mathbf{FB}^{\mathrm{op}}}\mathsf{triv}\quad\oplus\quad H(s)$$\cong$ $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),P^{\mathbf{FA}}_{\bullet})\ \ \cong\ \ \mathsf{sgn}_{s}\otimes_{\mathfrak{S}_{s}}\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{s}).$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),(\overline{P}^{\otimes*}/\Lambda)\odot_{\mathbf{FB}^{\mathrm{op}}}\mathsf{triv})\cong\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)\odot_{\mathbf{FB}^{\mathrm{op}}}\mathsf{triv}\cong\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),P^{\mathbf{FA}}_{\bullet}/\Lambda).$
Using the decomposition $P^{\mathbf{FA}}_{\bullet}\cong P^{\mathbf{FA}}_{\bullet}/\Lambda\ \oplus\ P^{\mathbf{FA}}_{\Lambda,\bullet}$, the first statement follows by combining Proposition 11.4 with Lemma 11.5. The statement for $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)]$ then follows by inverting $-\odot_{\mathbf{FB}^{\mathrm{op}}}\mathsf{triv}$ using Proposition 8.3; Proposition 8.9 is used to treat the contribution from the hook representations. Finally, the only difference when passing from $\overline{P}^{\otimes*}/\Lambda$ to $\overline{P}^{\otimes*}$ is a copy of $[\mathsf{sgn}_{s}]$ from the direct summand $\Lambda^{s}(\overline{P})$ of $\overline{P}^{\otimes s}$. The final statement follows from the equality $[\mathsf{sgn}_{s}]-\mathcal{S}(s)=\mathcal{S}(s+1)$ given by Lemma 8.7. ∎

### v3

@@ 10.1. Calculating $\mathrm{Hom}(\Lambda^{t}(P^{\mathbf{FA}}),P^{\mathbf{FA}}_{\mathbf{n}})$
Clearly $\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{t})=0$ if $s>t$, so we may assume that $s\leq t$. Now, as in the proof of Proposition 2.5, there is an isomorphism of $\mathbbm{k}(\mathfrak{S}_{t}\times\mathfrak{S}_{s}^{\mathrm{op}})$-modules $\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{t})\cong\mathbbm{k}\mathfrak{S}_{t}/\mathfrak{S}_{t-s}$, where the actions are inherited from the left and right regular actions on $\mathbbm{k}\mathfrak{S}_{t}$. It follows that one has the isomorphism $\mathsf{sgn}_{t}\otimes_{\mathfrak{S}_{t}}\mathbbm{k}{\mathbf{FI}}(\mathbf{s},\mathbf{t})\cong\mathsf{sgn}_{t}|_{\mathfrak{S}_{t-s}}\otimes_{\mathfrak{S}_{t-s}}\mathsf{triv}_{t-s}$ with respect to the residual $\mathbbm{k}\mathfrak{S}_{s}$-action arising from $\mathsf{sgn}_{t}|_{\mathfrak{S}_{s}}$ on the right hand side. This is clearly zero unless $t-s\in\{0,1\}$ in which case the representation identifies as stated. ∎
@@
This result is related to the short exact sequence given in Corollary 4.18.
@@
$0\rightarrow\mathbbm{k}_{\mathbf{0}}^{\mathbf{FA}^{\mathrm{op}}}\rightarrow\mathbbm{k}^{\mathbf{FA}^{\mathrm{op}}}\rightarrow\overline{\mathbbm{k}}^{\mathbf{FA}^{\mathrm{op}}}\rightarrow 0.$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},P^{\mathbf{FA}}_{\mathbf{n}})]=[\mathbbm{k}{\mathbf{FS}}(\mathbf{n},-)]\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(\mathbf{n},\mathbf{k}))\otimes[\mathsf{sgn}_{k-1}].$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},P^{\mathbf{FA}}_{\mathbf{n}})]=[\overline{P^{\mathbf{FA}}_{\mathbf{n}}}]\odot_{\mathbf{FB}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}P^{\mathbf{FA}}_{n}(\mathbf{k}))\otimes\mathcal{S}(k-1),$
The hypothesis $n>0$ ensures that $\overline{P^{\mathbf{FA}}_{\mathbf{n}}}=P^{\mathbf{FA}}_{\mathbf{n}}$. Thus, using the identity as $\mathbbm{k}\mathbf{FB}$-bimodules $\mathbbm{k}\mathbf{FA}\cong\mathsf{triv}\odot_{\mathbf{FB}}\mathbbm{k}{\mathbf{FS}}\cong\mathbbm{k}{\mathbf{FS}}\odot_{\mathbf{FB}}\mathsf{triv}$ given in Example 2.6, together with Proposition 8.3, one has $\overline{P^{\mathbf{FA}}_{\mathbf{n}}}]\odot_{\mathbf{FB}}\mathcal{S}(0)=[\mathbbm{k}{\mathbf{FS}}(\mathbf{n},-)]$, which accounts for the first term in the statement.
@@
$\sum_{k\geq 1}(\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}P^{\mathbf{FA}}_{n}(\mathbf{k}))\otimes\mathcal{S}(k-1)$ $=$ $\sum_{k\geq 1}M(\mathbf{n},\mathbf{k})\otimes(\mathcal{S}(k-1)\oplus\mathcal{S}(k))$
@@
where the first equality reindexes the sum. The result follows. ∎
@@
Hence, using Proposition 8.3, one can calculate $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]$:
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]$ $=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)$
$+\ \sum_{k\geq 1}\Big{(}[\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(-,\mathbf{k})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\Big{)}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}},$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]=[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P^{\mathbf{FA}}_{*}})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0).$
The result follows by using the identification of $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P^{\mathbf{FA}}_{*}})]$ given by Theorem 10.6. ∎
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\mathbb{P}_{*})]$ $=$ $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]\ +\ \sum_{\ell\geq 0}[\mathsf{sgn}_{\ell}]_{\mathbf{FB}^{\mathrm{op}}}\boxtimes[\mathsf{sgn}_{\ell+1}]_{\mathbf{FB}}$
$=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)$
$+\quad\sum_{k\geq 1}\big{(}[\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(-,\mathbf{k})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\big{)}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$
@@ 11. Calculating $\mathrm{Hom}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})$
The purpose of this section is to calculate the underlying $\mathbbm{k}\mathbf{FB}$-bimodule of
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*}).$
Theorem 10.8 described $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]$ in the Grothendieck group of $\mathbbm{k}\mathbf{FB}$-bimodules. We exploit the short exact sequence
$0\rightarrow\Lambda^{\bullet+1}(\overline{P})\rightarrow\mathbb{P}_{\bullet}\rightarrow\overline{P}^{\otimes\bullet}\rightarrow 0.$
@@ 11.1. Morphisms out of $\Lambda^{s}(\overline{P})$
using the $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-module structure of Proposition 2.11 on $\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{s})$. Moreover, with respect to this isomorphism, for $s,t\in\mathbb{N}$, the short exact sequence $0\rightarrow\Lambda^{s+1}(\overline{P})\rightarrow\Lambda^{s+1}(P^{\mathbf{FA}})\rightarrow\Lambda^{s}(\overline{P})\rightarrow 0$ induces a short exact sequence (writing $\mathrm{Hom}$ for $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}$)
$0\rightarrow\mathrm{Hom}(\Lambda^{s}(\overline{P}),P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow\mathrm{Hom}(\Lambda^{s+1}(P^{\mathbf{FA}}),P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow\mathrm{Hom}(\Lambda^{s+1}(\overline{P}),P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow 0$
@@
The first statement is an immediate consequence of the natural isomorphism
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes s},P^{\mathbf{FA}}_{\bullet})\cong\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{s})$
provided by Theorem 4.23. Moreover, Corollary 4.25 implies that the inclusion $\Lambda^{s+1}(\overline{P})\hookrightarrow\Lambda^{s+1}(P^{\mathbf{FA}}_{\mathbf{1}})$ induces a surjection
@@ 11.2. Calculating $\mathrm{Hom}(\Lambda^{s}(\overline{P}),\overline{P}^{\otimes*})$
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),\overline{P}^{\otimes\bullet}/\Lambda)\odot_{\mathbf{FB}^{\mathrm{op}}}\mathsf{triv}\ \oplus\ H(s)$ $\cong$ $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),P^{\mathbf{FA}}_{\bullet})$
$\cong$ $\mathsf{sgn}_{s}\otimes_{\mathfrak{S}_{s}}\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{s}).$
@@
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),(\overline{P}^{\otimes*}/\Lambda)\odot_{\mathbf{FB}^{\mathrm{op}}}\mathsf{triv})$ $\cong$ $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)\odot_{\mathbf{FB}^{\mathrm{op}}}\mathsf{triv}$
$\cong$ $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),P^{\mathbf{FA}}_{\bullet}/\Lambda).$
Using the decomposition $P^{\mathbf{FA}}_{\bullet}\cong P^{\mathbf{FA}}_{\bullet}/\Lambda\ \oplus\ P^{\mathbf{FA}}_{\Lambda,\bullet}$, the first statement follows by combining Proposition 11.4 with Lemma 11.5. The statement for $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{s}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)]$ then follows by inverting $\odot_{\mathbf{FB}^{\mathrm{op}}}\mathsf{triv}$ using Proposition 8.3; Proposition 8.9 is used to treat the contribution from the hook representations. Finally, the only difference when passing from $\overline{P}^{\otimes*}/\Lambda$ to $\overline{P}^{\otimes*}$ is a copy of $\mathsf{sgn}_{s}]$ from the direct summand $\Lambda^{s}(\overline{P})$ of $\overline{P}^{\otimes s}$. The final statement follows from the equality $\mathsf{sgn}_{s}]-\mathcal{S}(s)=\mathcal{S}(s+1)$ given by Lemma 8.7. ∎